If your consumption increases from $60,000/yr to $70,000/yr when your disposable income increases from $85,000 to $98,500/yr, calculate your MPC. *




1
Expert's answer
2021-10-19T15:59:01-0400

"MPC=\frac{Change in Consumer Spending}{Change in Income}"

"=\frac{70,000-60,000}{98,500-85,000}"

"=\frac{10,000}{13,500}"

"=0.74"